Find the general solution of the following differential equation: (1) ?′′ + 5?′ + 6? = 2????*?^? (2) ?′′ + 2?′ + ? = ? + ?e^(-t).

(please solve Question No.7 only)

7. (30 points) Find the general solution of the following differential equation: (1) y + 5y + 6y = 2etsint (2) y + 2y + y=
Answer #1

dy + by = 2 et sint lyr syt by - Ret sant writing of as we have Dy + 5Dy & by: 2 et sint. - 10+50+6) y 2 et sint which is faPIJ FODE et - e P. I = et (2 sint) D41 +2D + 5D +5+6 - (2 sint). D47D+12 Now, P.I. = sin at is given by, VP-1 - sinat f(D). I(2) g + 2y ty iedy + 2 dy dt dt ty=trte t+tet. => 01 4+te+ complašrentary function, Cfl (6+atem =t+tet e-t Writing of as D,PI. Joc tet tet 0+ 1)2 t since pa ps of fim eatv (OH)2 eat 2004) where is a function of t et Ilt), D D -t means +14) [6 inte